# EquiTabPFN: A Target-Permutation Equivariant Prior Fitted Network

### v2

@@ 6.2 Results
EquiTabPFN enables in-context-learning on datasets with unseen class counts. Figure 4 shows accuracies relative to the KNN baseline for all models on $76$ datasets with less than $10$ classes (left) and 10 datasets with more than $10$ classes (right). When considering datasets with more than $10$ classes (right), EquiTabPFN obtains the best median relative accuracy across all datasets. It strongly outperforms TabPFNv2, which performs worse than a linear model or random forests in terms of relative median. These results show that target-equivariance allows EquiTabPFN to seamlessly generalize to larger numbers of classes even though it was trained on datasets with less than $10$ classes, just like TabPFNv2. Additional evaluations in Tables 3 and 4 of Appendix C show that such improvement is consistent over various metrics: AUC, Accuracy and F1 score, for datasets with unseen class counts.
@@
Competitive performance on datasets with class counts seen during pre-training. On the 76 datasets with less than 10 classes (Figure 4, left), EquiTabPFN performs comparably to the state-of the art method, TabPFNv2, even though it did not benefit from the improved data prior used in Hollmann et al. [10] as it is not publicly available. To assess the impact of the pre-training prior on performance, we pre-trained the same network architecture as TabPFNv2 using the same publicly available protocol used for training our method and reported the results for reference (TabPFNv2). The results show that EquiTabPFN consistently outperforms TabPFNv2 both on dataset with less than $10$ classes and those with unseen class count. These results suggest that EquiTabPFN would likely benefit from the improved training procedure and prior of TabPFNv2.
@@
Speedup over TabPFNv2 on datasets with unseen class counts. As discussed in Section 3, PFNs models cannot natively handle an arbitrary number of classes due to its decoder architecture. In order to apply TabPFNv2 to problems with more than $10$ classes, we employ the error-correcting output codes (ECOC) strategy [4] as recommended in Hollmann et al. [10]. Such approach requires decomposing the classification problem into several smaller classification tasks with up to $10$ classes, performing predictions for each sub-task using an ensemble of TabPFNv2 models, at least one model per task, then aggregating predictions using an ECOC strategy. This incurs an extra computational cost for performing ensembling. All results reported in Figure 4 (right) are using the minimal number of ensembles to guarantee coverages of all classes. Despite this sophisticated aggregation strategy, the performance of TabPFNv2 degrades significantly, while still incurring a substantial slow-down compared to EquiTabPFN as shown by the average run-times reported by color in Figure 4 (left). Note that, for datasets with less than $10$ classes (Figure 4, left) ensembling is not required anymore by TabPFNv2 and the runtime of both methods are comparable.
@@
EquiTabPFN achieves the best tradeoff between performance and cost on datasets with unseen class counts. To further illustrate the improved trade-off between runtime and accuracy of our method, we show in Figure 5 (left) the improvement in AUC relatively to KNN for all methods when both TabPFNv2 and EquiTabPFN are allowed to have more ensembles. Ensembling generally improves performance of PFN models, as it allows to make more robust predictions by averaging them over transformed version of the datasets, ex: by applying permutations of the labels order as discussed in Section 5.2. While ensembling helps improve performance of TabPFNv2, the improvement is marginal compared to EquiTabPFN without ensembling and comes at a considerable computational cost, as also shown in Figure 5 (right) and on the critical difference diagrams in Figures 8 and 7 of Appendix C. Therefore, amongst top performing methods, EquiTabPFN achieves the best trade-off.
@@ 7 Conclusion
Limitations. EquiTabPFN requires a quadratic extra-cost with the number of target dimensions to perform self-attention. While many tabular problems have a relatively small number of target dimensions, this may become a problem for very large number of dimension and future work could consider efficient self-attention to address this issue.
Code for training and evaluating our model, will be released to enable reproducibility of our results.
@@ Appendix A Proofs
Proof of Proposition 5.3.
By Assumption (B) we can express $\mathcal{L}$ as an expectation of the form:
@@
By Assumption (A), $\ell$ is invariant to permutations, which allows to further write:
@@
We will show that $\mathcal{E}^{\text{equi}}(f)=\mathcal{L}(f)-\mathcal{L}({f}^{\text{equi}})$ is non-negative and vanishes only when $f$ is equivariant. This is a direct consequence of Jensen’s inequality applied to the strictly convex function $y\mapsto\ell(y,y^{\star})$ (Assumption (A)). Indeed, for any samples $(X,Y,X^{\star},Y^{\star})$, the following holds:
@@
where the first line follows by definition of ${f}^{\text{equi}}$ while the second line uses Jensen’s inequality. Further taking the expectation w.r.t. $p$ shows that $\mathcal{E}^{\text{equi}}(f)\geq 0$. If $\mathcal{E}^{\text{equi}}(f)=0$, then by the above inequality it holds that $\ell\left({f}^{\text{equi}}_{X,Y}\left(X^{\star}\right),Y^{\star}\right)=\ell\left(f_{X,Y}\left(X^{\star}\right),Y^{\star}\right)$ almost surely. However, since $\ell$ is strictly convex in its first argument (Assumption (A)), the previous equality is only possible when ${f}^{\text{equi}}=f$ almost surely, meaning that $f$ is equivariant.
@@
Here, the cross-product term equals 00 since $\mathbb{E}_{\sigma}\left[\sigma^{-1}\left(f_{X,\sigma(Y)}\left(X^{\star}\right)\right)\right]={f}^{\text{equi}}_{X,Y}\left(X^{\star}\right)$ by definition of ${f}^{\text{equi}}$. Hence, we have shown that:
@@ Appendix B Additional Experimental details
Architecture details. We use an EquiTabPFN network with $12$ self-attention layers alternating between both type of attention introduced in Section 4: 6 blocks $\text{SelfAtt}_{c}$ and 6 blocks $\text{SelfAtt}_{b}$. Each self-attention layer consists of a multi-head attention blocks with $4$ heads, embeddings of dimension $512$, and hidden layers of dimension $1024$. This choice ensures a fair comparison with the models used in Hollmann et al. [9], Müller et al. [17], since the number of parameters ($25.17M$) are of the same order when counting them as proposed in Kaplan et al. [11].
Datasets. For the datasets with less than $10$ classes, we collect the ones with less than $3000$ samples and $100$ features and retain the datasets that contain the same number of classes accross all folds and splits. For the datasets with more than $10$ classes, we filter in addition the ones able to run inference on an 80GB A100 GPU. The datasets obtained are given in Table 1 and Table 3.
@@ Appendix C Additional experiment results
C.1 Binary classification decision boundary
In Figure 6, we show the decision boundary on 3 binary classification datasets for multiple baselines. To illustrate the stability of the method, we do not do ensembling for TabPFN and EquiTabPFN.
@@ C.1 Binary classification decision boundary
C.2 Critical difference diagrams and performance metric tables
We show the critical diagram using Autorank implementation [8] in Figure 8 for datasets with more than 10 classes and Figure 7 for datasets with less than 10 classes. Critical diagrams show the average rank of each method (lower is better) and use a horizontal bar to show the methods statistically tied. We set the confidence level to $0.05$ and use default hyper-parameters while forcing non-parametric mode to ensure stability.
We also give aggregate results for datasets with more than 10 classes in Table 3 and less than 10 classes in Table 4.
Figure 7: Critical diagram on the 76 real-world datasets with less than $10$ classes from Table 1.
Figure 8: Critical diagram on the 10 real-world datasets with more than $10$ classes from Table 2.
[…]
@@ C.2 Critical difference diagrams and performance metric tables
Table 3: Aggregate accuracy, AUC, F1 and runtime for all methods on datasets with more than 10 classes. Results are ordered by the median relative accuracy improvement w.r.t KNN over the 10 datasets after averaging over the 10 different splits. Mean accuracies, AUC and F1 score are averaged over all splits and datasets. Numbers after the symbol +/+/-+ / - refer to the standard error of the mean over all splits and datasets.
[…]
@@
Table 4: Aggregate accuracy, AUC, F1 and runtime for all methods on datasets with less than 10 classes. Results are ordered by the median relative accuracy improvement w.r.t KNN over the 10 datasets after averaging over the 10 different splits. Mean accuracies, AUC and F1 score are averaged over all splits and datasets. Numbers after the symbol +/+/-+ / - refer to the standard error of the mean over all splits and datasets.

### v4

@@ 6.2 Main results
EquiTabPFN enables in-context-learning on datasets with unseen class counts. Figure˜4 shows accuracies relative to the KNN baseline for all models on $76$ datasets with less than $10$ classes (left) and 10 datasets with more than $10$ classes (right). When considering datasets with more than $10$ classes (right), EquiTabPFN obtains the best median relative accuracy across all datasets. It strongly outperforms TabPFNv2, which performs worse than a linear model or random forests in terms of relative median. These results show that target-equivariance allows EquiTabPFN to seamlessly generalize to larger numbers of classes even though it was trained on datasets with less than $10$ classes, just like TabPFNv2. Additional evaluations in Tables˜5 and 6 of Appendix˜C show that such improvement is consistent over various metrics: AUC, Accuracy and F1 score, for datasets with unseen class counts.
@@
Competitive performance on datasets with class counts seen during pre-training. On the 76 datasets with less than 10 classes (Figure˜4, left), EquiTabPFN performs comparably to the state-of the art method, TabPFNv2, even though it did not benefit from the improved data prior used in Hollmann et al. [13] as it is not publicly available. To assess the impact of the pre-training prior on performance, we pre-trained the same network architecture as TabPFNv2 using the same publicly available protocol used for training our method and reported the results for reference (TabPFNv2). The results show that EquiTabPFN consistently outperforms TabPFNv2 both on dataset with less than $10$ classes and those with unseen class count. These results suggest that EquiTabPFN would likely benefit from the improved training procedure and prior of TabPFNv2.
@@
Speedup over TabPFNv2 on datasets with unseen class counts. As discussed in Section˜3, PFNs models cannot natively handle an arbitrary number of classes due to its decoder architecture. In order to apply TabPFNv2 to problems with more than $10$ classes, we employ the error-correcting output codes (ECOC) strategy [5] as recommended in Hollmann et al. [13]. Such approach requires decomposing the classification problem into several smaller classification tasks with up to $10$ classes, performing predictions for each sub-task using an ensemble of TabPFNv2 models, at least one model per task, then aggregating predictions using an ECOC strategy. This incurs an extra computational cost for performing ensembling. All results reported in Figure˜4 (right) are using the minimal number of ensembles to guarantee coverages of all classes. Despite this sophisticated aggregation strategy, the performance of TabPFNv2 degrades significantly, while still incurring a substantial slow-down compared to EquiTabPFN as shown by the average run-times reported by color in Figure˜4 (left). Note that, for datasets with less than $10$ classes (Figure˜4, left) ensembling is not required anymore by TabPFNv2 and the runtime of both methods are comparable.
@@
EquiTabPFN achieves the best tradeoff between performance and cost on datasets with unseen class counts. To further illustrate the improved trade-off between runtime and accuracy of our method, we show in Figure˜5 (left) the improvement in AUC relatively to KNN for all methods when both TabPFNv2 and EquiTabPFN are allowed to have more ensembles. Ensembling generally improves performance of PFN models, as it allows to make more robust predictions by averaging them over transformed version of the datasets, ex: by applying permutations of the labels order as discussed in Section˜5.2. While ensembling helps improve performance of TabPFNv2, the improvement is marginal compared to EquiTabPFN without ensembling and comes at a considerable computational cost, as also shown in Figure˜5 (right) and on the critical difference diagrams in Figures˜8 and 7 of Appendix˜C. Therefore, amongst top performing methods, EquiTabPFN achieves the best trade-off.
@@ 7 Conclusion
Non-equivariance. In some cases, the data may not be target equivariant, for instance on ordinal data. We found only a few of those cases in the benchmarks used (5 of the 86 datasets surveyed). Handling such cases would require to update the prior as it is currently equivariant to target permutation. The method could then be adapted by providing positional embedding or using column indices as input.
Limitations. EquiTabPFN requires a quadratic extra-cost with the number of target dimensions to perform self-attention. While many tabular problems have a relatively small number of target dimensions, this may become a problem for a very large number of dimensions and future work could consider efficient self-attention to address this issue. This may be alleviated by work focused on improving the efficiency of PFNs models [28, 38].
The code for training and evaluating our model is available at https://github.com/MichaelArbel/EquiTabPFN/.
@@ Appendix A Proofs
Proof of Proposition˜5.3.
By Assumption˜(B) we can express $\mathcal{L}$ as an expectation of the form:
@@
By Assumption˜(A), $\ell$ is invariant to permutations, which allows to further write:
@@
We will show that $\mathcal{E}^{\text{equi}}(f)=\mathcal{L}(f)-\mathcal{L}({f}^{\text{equi}})$ is non-negative and vanishes only when $f$ is equivariant. This is a direct consequence of Jensen’s inequality applied to the strictly convex function $y\mapsto\ell(y,y^{\star})$ (Assumption˜(A)). Indeed, for any samples $(X,Y,X^{\star},Y^{\star})$, the following holds:
@@
where the first line follows by definition of ${f}^{\text{equi}}$ while the second line uses Jensen’s inequality. Further taking the expectation w.r.t. $p$ shows that $\mathcal{E}^{\text{equi}}(f)\geq 0$. If $\mathcal{E}^{\text{equi}}(f)=0$, then by the above inequality it holds that $\ell\left({f}^{\text{equi}}_{X,Y}\left(X^{\star}\right),Y^{\star}\right)=\ell\left(f_{X,Y}\left(X^{\star}\right),Y^{\star}\right)$ almost surely. However, since $\ell$ is strictly convex in its first argument (Assumption˜(A)), the previous equality is only possible when ${f}^{\text{equi}}=f$ almost surely, meaning that $f$ is equivariant.
@@
Here, the cross-product term equals 0 since $\mathbb{E}_{\sigma}\left[\sigma^{-1}\left(f_{X,\sigma(Y)}\left(X^{\star}\right)\right)\right]={f}^{\text{equi}}_{X,Y}\left(X^{\star}\right)$ by definition of ${f}^{\text{equi}}$. Hence, we have shown that:
@@ Appendix B Additional Experimental details
Architecture details. We use an EquiTabPFN network with $12$ self-attention layers alternating between both type of attention introduced in Section˜4: 6 blocks $\text{SelfAtt}_{c}$ and 6 blocks $\text{SelfAtt}_{b}$. Each self-attention layer consists of a multi-head attention blocks with $4$ heads, embeddings of dimension $512$, and hidden layers of dimension $1024$. This choice ensures a fair comparison with the models used in Hollmann et al. [12], Müller et al. [22], since the number of parameters ($25.17M$) are of the same order when counting them as proposed in Kaplan et al. [14].
Datasets. For the datasets with less than $10$ classes, we collect the ones with less than $3000$ samples and $100$ features and retain the datasets that contain the same number of classes accross all folds and splits. For the datasets with more than $10$ classes, we filter in addition the ones able to run inference on an 80GB A100 GPU. The datasets obtained are given in Table˜1 and Table˜5.
@@ Appendix C Additional experiment results
C.1 Ablation on the different components of EquiTabPFN.
Table˜3 reports relative error reduction over TabPFN in two settings: (1) TabPFN bb + Eq dec. using the TabPFN backbone (without bi-attention) and our equivariant decoder, and (2) Bi-attn bb + MLP dec. using a bi-attention backbone and a standard MLP decoder from TabPFNv1. Since the encoder is a simple linear embedding, it is considered part of the backbone: fully connected for TabPFN-style models and 1x1 convolution for biattention-based ones. The combination of both—biattention and the equivariant decoder, e.g., the EquiTabPFN model we propose—yields the largest performance gain. We also tried using TabPFNv2 architecture backbone and modified it to make it target equivariant, then trained it using the publicly available code for the prior used for training TabPFNv1. This led to improvements compared to TabPFNv2 (the version we retrained ourselves using publicly available training prior). However, we found that using TabPFNv1’s backbone yielded the best performance overall.
EquiTabPFN TabPFN bb + Eq dec. Bi-attn bb + MLP dec.
% Error reduction +1.50% +0.94% -0.12%
Table 3: Error reduction over TabPFN for different model configurations.
C.2 Computational cost comparisons
Run time comparison. Table˜4 shows the run time comparison between EquiTabPFN, TabPFNv1 and TabPFNv2 on both types of datasets (small or large number of classes). On a small number of classes, EquiTabPFN incurs a slowdown of 5x compared to TabPFNv1. This is expected as TabPFNv1 was optimized to handle data with less than 10 classes. On datasets with more than 10 classes, the gap narrows (only 1.3x slowdown) as TabPFNv1 requires ensembling techniques to handle the larger number of classes. A more complete picture accounts for the tradeoff between time cost and performance as in Figure˜5 (left) and shows a clear advantage of EquiTabPFN in terms of efficiency.
EquiTabPFN TabPFNv1 TabPFNv2
< 10 Classes 0.12 0.02 0.16
> 10 Classes 0.40 0.30 2.80
Table 4: Time comparison (in seconds) across two types of datasets, depending on their class count.
FLOPS comparaison. While EquiTabPFN and TabPFN have similar parameter counts, EquiTabPFN uses more FLOPS. On an A100 GPU with 2,000 samples (100 features, 10 classes), EquiTabPFN required 566 GFLOPS vs. 76 GFLOPS for TabPFN ($\sim$7.45× more). However, with more classes (15), the gap narrows due to the ensembling needed for TabPFN to handle more than 10 classes (EquiTabPFN: 820 GFLOPS; TabPFN: 456 GFLOPS), consistent with runtime trends. Overall, those numbers remain small for a modern GPU given that a single H100 can easily reach 400 TFLOPS on an LLM training workflow for instance.
Memory cost. EquiTabPFN incurs an increase compared to TabPFN which was translated in the use of smaller batch size (first dimension of the activation tensors). However, the context (in terms of the number of samples that can be processed by the model on our devices) was not affected in the experiments. This is likely due to the moderate size of the contexts used whose ranges are within the recommended limits for TabPFN.
C.3 Binary classification decision boundary
In Figure˜6, we show the decision boundary on 3 binary classification datasets for multiple baselines. To illustrate the stability of the method, we do not do ensembling for TabPFN and EquiTabPFN.
@@ C.3 Binary classification decision boundary
C.4 Critical difference diagrams and performance metric tables
We show the critical diagram using Autorank implementation [11] in Figure˜8 for datasets with more than 10 classes and Figure˜7 for datasets with less than 10 classes. Critical diagrams show the average rank of each method (lower is better) and use a horizontal bar to show the methods statistically tied. We set the confidence level to $0.05$ and use default hyper-parameters while forcing non-parametric mode to ensure stability.
We also give aggregate results for datasets with more than 10 classes in Table˜5 and less than 10 classes in Table˜6.
Figure 7: Critical diagram on the 76 real-world datasets with less than $10$ classes from Table˜1.
Figure 8: Critical diagram on the 10 real-world datasets with more than $10$ classes from Table˜2.
Median relative Acc. Mean Acc. Mean AUC Mean F1 Mean time (s)
@@ C.4 Critical difference diagrams and performance metric tables
Table 5: Aggregate accuracy, AUC, F1 and runtime for all methods on datasets with more than 10 classes. Results are ordered by the median relative accuracy improvement w.r.t KNN over the 10 datasets after averaging over the 10 different splits. Mean accuracies, AUC and F1 score are averaged over all splits and datasets. Numbers after the symbol +/+/- refer to the standard error of the mean over all splits and datasets.
Median relative Acc. Mean Acc. Mean AUC Mean F1 Mean time (s)
@@
Table 6: Aggregate accuracy, AUC, F1 and runtime for all methods on datasets with less than 10 classes. Results are ordered by the median relative accuracy improvement w.r.t KNN over the 10 datasets after averaging over the 10 different splits. Mean accuracies, AUC and F1 score are averaged over all splits and datasets. Numbers after the symbol +/+/- refer to the standard error of the mean over all splits and datasets.